# Differentially Private Matchings

### v1

@@
Motivated by the importance of computing matchings in sensitive graph data, we initiate the study of differentially private algorithms for computing maximal and maximum matchings in general graphs. We provide a number of algorithms and lower bounds for this problem in different models and settings.
@@
  • We then consider implicit solutions, where at the end of the computation there is an ($\varepsilon$-differentially private) billboard and each node can determine its matched edge(s) based on what is written on this publicly visible billboard. For this solution concept, we provide tight upper and lower (bicriteria) bounds, where the degree bound is violated by a logarithmic factor (which we show is necessary).
  • We further show that our algorithm can be made distributed in the local edge DP (LEDP) model, and can even be done in a logarithmic number of rounds if we further relax the degree bounds by logarithmic factors.
  • Our edge-DP matching algorithms give rise to new matching algorithms in the node-DP setting by combining our edge-DP algorithms with a novel use of arboricity sparsifiers. Interestingly, we prove an impossibility result for publicly releasing such sparsifiers under differential privacy, even though they form a key component of our node-DP algorithm. Our techniques also allow us to improve the bipartite results of [hsu2014private] by a polylogarithmic factor.
  • Finally, we demonstrate that all of our results can also be implemented in the continual release model.
@@ 1 Introduction
Differential privacy [DMNS06] is the gold standard for protecting the privacy of individuals. While differential privacy was originally introduced in the context of numerical and statistical databases, nothing in its definition requires this. And since many problems in graph algorithms are actually problems about graph analytics over graphs containing sensitive personal information, there has been a large amount of important work on differentially private graph algorithms. This includes differentially private versions of graph problems like triangle counting, shortest paths, graph cuts, densest subgraph, $k$-core decompositions, and many others [NRS07, GuptaLMRT10, BGM22, DMN23, DLRSSY22, ELRS22, kasiviswanathan2013analyzing, kalemaj2023node, LUZ24, mueller2022sok, RS16, raskhodnikova2016differentially, Upadhyay13]. In the graph setting, differential privacy is generally defined in terms of edge-neighboring graphs or node-neighboring graphs where the graphs differ in one edge or all edges incident to any one node, respectively; this intuitively corresponds to requiring that when we run our algorithm on neighboring graphs, the distribution of outputs is very similar (see Section 2 for formal definitions).
@@
  1. 1.
    It is limited to bipartite graphs, since that is the usual setting for allocations (where one side consists of agents and the other side consists of goods and a matching indicates which agents get which goods). However, there are obviously many settings in which we might want a private matching (or $b$-matching) in a general graph. For example, consider online dating where a matching should be kept private but individuals in a pair should know who they are matched with. Alternatively, consider conference reviewing, where people might be both authors and reviewers and the edges between reviewers and authors should be private.
  2. 2.
    It only considers the setting of node privacy, since in an allocation problem the valuation function for each agent is sensitive information. However, in many settings the natural notion of privacy is edge-based; for example, in social network analysis, we may wish to compute matchings without revealing information about the relationship between two specific nodes. In such settings, while node privacy suffices, it is not necessary. As node privacy is a stronger notion, there are often prohibitively strong lower bounds against it. So in these settings, it would be impractical to depend on node private algorithms (and suffer their lower bounds) when edge privacy would be sufficient and would allow for better guarantees. This is a fundamental reason for the explosion of recent interest in edge private algorithms.
  3. 3.
    It gives a joint differentially private approximate maximum matching algorithm. Informally, joint differential privacy [KPRU14] is a weaker definition of privacy where for every node $i\in V$, privacy is maintained when the solution output of $i$ is not considered in the set of outputs. This definition is still strong in the sense that private information can only influence the output of any one node but not the composite outputs of many nodes; however, it is a much weaker definition than differential privacy. While joint differential privacy may make sense in an allocation setting (see [hsu2014private] for more discussion), it is not clear that we need to relax differential privacy all the way to joint differential privacy if we care about matchings rather than allocations.
Thus, we still have a very natural question: is it possible to give differentially private algorithms for maximum (or maximal) matchings in general graphs, i.e., for settings other than allocations? In this paper we introduce the study of privately computing matchings in general graphs. We give the first algorithms and lower bounds for this question in a variety of models including both edge- and node-DP, local privacy, and continual release. As a corollary of our techniques, we improve the bounds of [hsu2014private] for the node-private joint DP bipartite setting.
@@ 1.1 Our Contributions
We now state our results in the various settings mentioned above.
@@ 1.1.1 Solution Types
It is easy to see that given an input graph, it is impossible to differentially privately output any subgraph (including a matching) containing only true edges; doing so would obviously reveal the existence of any edge that we output. So we must turn to other ways to output a matching that will not result in privacy violations. This is one reason why [hsu2014private] resorted to joint differential privacy.
@@
The proof of Theorem 1.1 is presented in Section 3. To interpret this theorem, think of $\delta\leq 1/n$ (so the $\delta$ term does not contribute anything meaningful) and small constant $\varepsilon$. Then a few meaningful regimes of $b,\gamma,\beta$ include the following.
@@
Theorem 1.1 suggests that we must find another way to return a subset of edges representing a matching. A natural approach, which we adopt through the rest of this paper, is to use “implicit” solutions in the sense of [GuptaLMRT10], which was later formalized by [hsu2014private] as the billboard model. In this model there is a billboard of public information released by the coordinator or the nodes themselves. Using the billboard of public information and its own privately stored information, each node can privately determine which of its incident edges are in the solution (if any). So in this model, we output something (to the billboard) that is “locally decodable” at every node (where each node can also use its private information in the decoding). Henceforth, we refer to the information on the billboard as an “implicit solution”. We formally define this implicit solution representation in Definition 2.4. It is important to note that the billboard model is not a privacy model, but a solution release model.
@@ 1.1.2 Edge Privacy (LEDP)
Our results for edge privacy combine the billboard model of implicit solutions with a strong local version of privacy known as local edge differential privacy (LEDP) [IMC21locally, DLRSSY22]. In the LEDP model, nodes do not reveal their private information to anyone. Rather, nodes communicate with the curator over (possibly many) rounds, where in each round each node releases differentially private outputs which are accumulated in a global transcript. The transcript is publicly visible and must also be differentially private. It is very natural to combine the LEDP and billboard models by thinking of the billboard as this differentially private transcript, which is the approach we take.
We note that although both [hsu2014private] and our results are in the billboard model for solution release, we use LEDP rather than centralized differential privacy as our privacy model, and hence our solutions are in a much stronger privacy model (although only for edge-privacy rather than for node-privacy).
As with our lower bound in Theorem 1.1, we will allow our algorithms to be bicriteria and output $b$-matchings. Our solutions are additionally guaranteed to contain (approximate) maximum matchings, with high probability. Since our main goal is to privately output matchings rather than $b$-matchings, it makes sense to measure the quality of the solution using the size of the largest contained matching rather than the size of the $b$-matching. Indeed, consider the graph consisting of the disjoint union of a matching of size $\log(n)$ and a star graph on $\log(n)$ vertices. All other vertices are isolated vertices. Then for $b=\log(n)$, outputting the matching or the star would both be a $\nicefrac{{1}}{{2}}$-approximate $b$-matching. However, the latter is qualitatively useless as an approximate matching!
@@
Theorem 1.2 (LEDP Maximal Matching).
For $\varepsilon\in(0,1)$, there is an $\varepsilon$-LEDP algorithm that, with high probability, outputs an (implicit) $b$-matching in the billboard model for $b=O(\log(n)/\varepsilon)$ that contains a maximal matching.
Since the size of a maximum matching is at most twice the size of any maximal matching, the above theorem immediately implies a $2$-approximate maximum matching algorithm.
Corollary 1.3.
There is an $\varepsilon$-LEDP algorithm that, with high probability, outputs an implicit $O(\log(n)/\varepsilon)$-matching in the billboard model which contains a $2$-approximate maximum matching.
@@
Theorem 1.4 (Efficient Maximal Matching).
For $\varepsilon\in(0,1)$, there is an $\varepsilon$-LEDP algorithm that terminates in $O(\log n)$ rounds and, with high probability, outputs an implicit $b$-matching in the billboard model for $b=O(\log^{2}(n)/\varepsilon)$ that contains a maximal matching.
Corollary 1.5.
There is an $\varepsilon$-LEDP algorithm that, with high probability, outputs an implicit $O(\log^{2}(n)/\varepsilon)$-matching in the billboard model which contains a $2$-approximate maximum matching.
Our algorithms not only use ideas from parallel and distributed algorithms but also use new formalizations of privacy tools which may be of independent interest outside of the scope of this paper, including the following three techniques: 1) our novel Public Vertex Subset Mechanism, 2) a formalization of the adaptive Laplace mechanism, and 3) new usages of concurrent composition within graph algorithms. Our Public Vertex Subset Mechanism allows us to differentially privately release a set of candidate vertices per node. Each node can then intersect the released set with their private adjacency list to produce implicit vertex subset solutions. Our mechanism is an independent contribution that may be useful for other implicit graph problems beyond matchings.
We complement our upper bounds with a lower bound which shows that for the particular type of implicit solutions used in Theorem 1.2, our reliance on $b$ is necessary: if $b$ is sublogarithmic then it cannot be a good approximation to the maximal matching. The following theorem is an informal version of this; see Section 9.1 and Theorem 9.1 for the precise statement.
@@
Let $\mathcal{A}$ be an algorithm which satisfies $\varepsilon$-edge DP and outputs an implicit solution of the type used in Theorem 1.2 which is a $(1+\eta)$-approximate maximal matching [^2] Given a graph $G=(V,E)$, a $b$-matching $F$ is said to be a $\psi$-approximate maximal matching if there exists a maximal matching with size $M$ such that $F$ has size at least $M/\psi$. If there is a maximal $b$-matching of size $M$ where $F$ has size at least $M/\psi$, then we say that $F$ is a $\psi$-approximate $b$-matching. with probability at least $1-1/\mathrm{poly}(n)$. Then, when decoded from the billboard, with probability at least $1/2$, there is a node matched with at least $\Omega\left(\frac{\log(n)}{\varepsilon}\right)$ other nodes.
@@ 1.1.3 Node Privacy
In the node differential privacy setting, two graphs are considered neighbors if one can be obtained from the other by replacing all edges incident on any one node. It is a significantly more challenging setting than edge privacy. We show in Section 8 that the ideas we developed for edge-DP in general graphs can be combined with a “matching sparsifier” [solomon2021local] to give the first implicit node-DP algorithms for matching in general graphs. Specifically, we show the following upper bound results in node-DP where $\alpha$ is the arboricity [^3] Recall that a graph has arboricity $\alpha$ if it can be partitioned into $\alpha$ edge-disjoint forests. of the input graph.
@@
Let $\eta\in(0,1]$, $\varepsilon\in(0,1)$, and $\alpha$ be the arboricity of the input graph. There is an $\varepsilon$-node DP algorithm that outputs an implicit $b$-matching. Moreover, with high probability, (i) $b=O\left(\frac{\alpha\log^{2}(n)}{\eta\varepsilon}\right)$, and (ii) the implicit solution contains a $\left(2+\eta,O\left(\frac{\log^{2}(n)}{\varepsilon}\right)\right)$-approximate maximum matching.
Our techniques also lead to improved results for the setting of [hsu2014private]: bipartite graphs with node differential privacy in the billboard model [^4] They show that this implies joint (node)-differential privacy, and all results are stated in terms of joint differential privacy.. In this model, they view the nodes in the “left side” of the bipartition as goods and the “right side” nodes of the graph as bidders for these goods. Two such bipartite graphs are neighboring if one can be obtained from the other by adding/deleting any one bidder (and all its incident edges).
[hsu2014private] show guarantees of the following form. Suppose each item has supply $s$, meaning that each item can be matched with $s$ bidders (let us call these $s$-matchings); if $s=\Omega(\log^{3.5}{n}/\varepsilon)$, then they give an algorithm which (implicitly) outputs a $(1+\eta)$-maximum matching while guaranteeing differential privacy. [^5] They also consider the weighted version of the problem and further generalizations where each bidder has gross substitutes valuations. Our techniques also extend to these cases with improved results, but we limit our discussion here to matchings as that is the focus of this work. They also show that $s=\Omega(1/\sqrt{\eta})$ is necessary to obtain any guarantees, leaving open the question of tightening these bounds. We substantially tighten the upper bound, showing that $s=\Theta(\log{n}/\varepsilon)$ suffices:
Theorem 1.8.
Let $\eta\in(0,1/2]$ and $\varepsilon\in(0,1)$. There is an $\varepsilon$-node DP algorithm for bipartite graphs which outputs a $(1+\eta)$-approximate maximum $s$-matching in the billboard model, for sufficiently large $s\geq\Omega(\log(n)/(\eta^{4}\varepsilon))$ with probability at least $1-\frac{1}{n^{c}}$ for constant $c\geq 1$.

### v2

@@
Motivated by the importance of computing matchings in sensitive graph data, we study differentially private algorithms for computing maximum matchings and $b$-matchings in general graphs. Our algorithms return implicit solutions for these problems. We provide a number of algorithms and lower bounds for these problems in different models and settings.
@@
  • We then consider implicit solutions, where at the end of the computation there is an ($\varepsilon$-differentially private) billboard and each node can determine its matched edge(s) based on what is written on this publicly visible billboard. For this solution concept, we provide tight upper and lower (bicriteria) bounds, where the degree bound is violated by a logarithmic factor, which we show is necessary via a novel lower bound.
  • We further show that our algorithm can be implemented in the local edge DP (LEDP) model, and can even be done in a logarithmic number of rounds if we further relax the degree bounds by logarithmic factors.
  • Moreover, we implement our algorithms in the node-DP setting by combining edge-DP algorithms with a novel use of arboricity sparsifiers. Interestingly, we prove an impossibility result for publicly releasing such sparsifiers under differential privacy, even though they form a key component of our node-DP algorithm. Our techniques also allow us to improve the bipartite results of [hsu2014private] by a polylogarithmic factor.
  • We demonstrate that all of our results can also be implemented in the continual release model under both edge- and node-privacy.
  • Finally, we extend our main algorithmic result to the $b$-matching setting, including an extension of our node-DP algorithm using a novel approximate $b$-matching arboricity sparsifier. Our sparsifier may be of independent interest even in the non-private graph setting.
@@ 1 Introduction
Differential privacy [DMNS06] is the gold standard for protecting the privacy of individuals. While differential privacy was originally introduced in the context of numerical and statistical databases, a growing number of works have placed its focus on sensitive graph data. Since many problems in graph algorithms are actually problems about graph analytics over graphs containing sensitive personal information, there has been a large amount of important work on differentially private graph algorithms. This includes differentially private versions of graph problems like triangle counting, shortest paths, graph cuts, densest subgraph, $k$-core decompositions, and many others [NRS07, GuptaLMRT10, BGM22, DMN23, DLRSSY22, ELRS22, kasiviswanathan2013analyzing, kalemaj2023node, LUZ24, mueller2022sok, RS16, raskhodnikova2016differentially, Upadhyay13]. In the graph setting, differential privacy is generally defined in terms of edge-neighboring graphs or node-neighboring graphs where the graphs differ in one edge or all edges incident to any one node, respectively; this intuitively corresponds to requiring that when we run our algorithm on neighboring graphs, the distribution of outputs is very similar (see Section 2 for formal definitions).
@@
Thus, very natural and important questions remain: is it possible to give differentially private algorithms for maximum (or maximal) matchings and $b$-matchings in general graphs beyond bipartite graphs? In this paper, we introduce the study of privately computing matchings and $b$-matchings in general graphs. We give the first algorithms and lower bounds for this question in a variety of models including both edge- and node-DP, local privacy, and continual release via a variety of novel techniques, some of which may be of independent interest in both the privacy and non-private graph algorithms communities. As a corollary of our techniques, we improve the bounds of [hsu2014private] for the node-private bipartite setting.
@@ 1.1 Our Contributions
We now state our results and our main technical contributions in the various settings mentioned above.
@@ 1.1.1 Solution Types
It is easy to see that given an input graph, it is impossible to differentially privately output any subgraph (including a matching) containing only true edges; doing so would obviously reveal the existence of any edge that we output. So we must turn to other ways to output a matching that will not result in privacy violations.
@@
The proof of Theorem 1.1 is presented in Section 3. To interpret this theorem, think of $\delta\leq 1/n$ (so the $\delta$ term does not contribute anything meaningful) and small constant $\varepsilon$. Then a few meaningful regimes of $b,\gamma,\beta$ include the following:
@@
We can actually strengthen this into a lower bound even on fractional matchings. A fractional $b$-matching is an assignment of values in $[0,1]$ to edges (i.e., a vector $X\in[0,1]^{\binom{V}{2}}$) so that $\sum_{v\in V}X_{\{u,v\}}\leq b$. As the proof of the following theorem proceeds similarly to to lower bound for explicit integral solutions, we defer it to Appendix B.
Theorem 1.2 (Lower Bound on Explicit Fractional Solutions).
Let $\mathcal{A}$ be an algorithm which satisfies $(\varepsilon,\delta)$-edge DP and always outputs a fractional $b$-matching $X_{G}^{\mathcal{A}}\in[0,1]^{\binom{V}{2}}$. If $\mathcal{A}$ outputs a $(\gamma,\beta)$-approximation (even just in expectation) of the maximum matching of a $0/1$-weighted (Definition 2.7) input graph $G=(V,E)$, then
$\gamma(e^{2\varepsilon}b+\delta n)+\beta\geq n/2.$
@@
Theorem 1.1 suggests that we must find another way to return a subset of edges representing a matching. A natural approach, which we adopt through the rest of this paper, is to use “implicit” solutions in the sense of [GuptaLMRT10], which was later formalized by [hsu2014private] as the billboard model. In this model there is a billboard of public information released by the coordinator or the nodes themselves. Using the billboard of public information and its own privately stored information, each node can privately determine which of its incident edges are in the solution (if any). So in this model, we output something (to the billboard) that is “locally decodable” at every node (where each node can also use its private information in the decoding). That is, the information released to the billboard is $\varepsilon$-differentially private and each node computes its own part of the solution by combining the public information on the billboard with its own private adjacency list. Henceforth, we refer to the information on the billboard as an “implicit solution”. We formally define this implicit solution representation in Definition 2.4; such a solution has been used previously in the context of vertex cover, set cover, and maximum matching in bipartite graphs [GuptaLMRT10, hsu2014private]. It is important to note that the billboard model is not a privacy model, but a solution release model.
@@ 1.1.2 Edge Privacy (LEDP)
Our results for edge privacy are within the novel setting that is a combination of the billboard model of implicit solutions with a strong local version of privacy known as local edge differential privacy (LEDP) [IMC21locally, DLRSSY22]. In the LEDP model, nodes do not reveal their private information to anyone. Rather, nodes communicate with the curator over (possibly many) rounds, where in each round each node releases differentially private outputs which are accumulated in a global transcript. The transcript is publicly visible and must also be differentially private. It is very natural to combine the LEDP and billboard models by thinking of the billboard as this differentially private transcript, which is the approach we take.
Although both [hsu2014private] and our results are in the billboard model for solution release, their algorithm hinges on the “binary mechanism” [CSS11], a mechanism for releasing a stream of $n$ private counting queries with only $O(\log n)$ error, rather than the $O(\operatorname{poly}(n))$ error attained through naive composition. This mechanism releases noisy partial sums such that each query can be answered by combining $O(\log n)$ partial sums. However, this requires a centralized curator who can compute the true partial sums before adding noise, which seems incompatible with the local privacy model.
In comparison, we can guarantee LEDP rather than centralized differential privacy, and hence our solutions are in a much stronger privacy model (although only for edge-privacy rather than for node-privacy). We also later provide solutions for node-DP in the centralized setting.
As with our lower bound in Theorem 1.1, we will allow our algorithms to be bicriteria and output matchings that have up to $b$-degree. Interestingly, we present three main results for matchings depending on the size of $b$.
Our solutions are additionally guaranteed to (approximate) maximum matchings, with high probability. Given a large enough gap between $b$ and $b^{\prime}$, we also show that our solutions are guaranteed to (approximate) maximum $b^{\prime}$-matchings. When $b$ is sufficiently large, our returned matching will be a $(2+\eta)$-approximation of the maximum $b$-matching, and would no longer need to be a bicriteria approximation.
@@
Theorem 1.3 (LEDP $2$-Approximate Maximum Matching).
For $\varepsilon\in(0,1)$, Algorithm 2 is an $\varepsilon$-LEDP algorithm that, with high probability, outputs an (implicit) matching with degree at most $b$, in the billboard model, for $b=O(\log(n)/\varepsilon)$ that has the size of a $2$-approximate maximum matching.
We further extend our result to $b^{\prime}$-matchings where given a sufficient gap between $b$ and $b^{\prime}$, we can obtain an $\varepsilon$-LEDP algorithm that outputs a matching with degree at most $b$ that is a $2$-approximate maximum $b^{\prime}$-matching. We present these results in Section 10.
Theorem 1.4 (LEDP $2$-Approximate Maximum $b^{\prime}$-Matching).
For $\varepsilon\in(0,1)$ and $\eta\in(0,1)$, Algorithm 9 is an $\varepsilon$-LEDP algorithm that, with high probability, outputs an (implicit) matching with degree at most $b$, in the billboard model, for $b\geq\frac{(1+\eta)^{2}}{1-\eta}\cdot b^{\prime}+O\left(\frac{\log(n)}{\eta^{2}\varepsilon}\right)$ that has the size of a $2$-approximate maximum $b^{\prime}$-matching.
Finally, when $b$ is sufficiently large, we no longer require a bicriteria approximation; instead, we are guaranteed to return an approximate maximum $b$-matching, with high probability. We present these results in detail in Section 10.
Theorem 1.5 (LEDP $2$-Approximate Maximum $b$-Matching).
For $\varepsilon\in(0,1)$ and $\eta\in(0,1/2)$, Algorithm 9 is an $\varepsilon$-LEDP algorithm that, with high probability, outputs an (implicit) $b$-matching, in the billboard model, for $b=\Omega\left(\frac{\ln(n)}{\eta^{3}\varepsilon}\right)$ that has the size of a $(2+\eta)$-approximate maximum $b$-matching.
@@
Theorem 1.6 (Efficient 2-Approximate Maximum Matching).
For $\varepsilon\in(0,1)$, Algorithm 3 is an $\varepsilon$-LEDP algorithm that terminates in $O(\log n)$ rounds and, with high probability, outputs an (implicit) matching with degree at most $b$ in the billboard model for $b=O(\log^{2}(n)/\varepsilon)$ that has the size of a $2$-approximate maximum matching.
We complement our upper bounds with a lower bound which shows that even for the particular type of implicit solutions used in Theorem 1.3, our reliance on $b$ is necessary: if $b$ is sublogarithmic, then even an implicit solution cannot be a good approximation to the maximum matching in the graph. The following theorem is an informal version of this; see Section 11.1 and Theorem 11.1 for the precise statement.
@@
Let $\mathcal{A}$ be an algorithm which satisfies $\varepsilon$-edge DP and outputs an implicit solution of the type used in Theorem 1.3 which is the size of a $\gamma$-approximate maximal matching [^2] Given a graph $G=(V,E)$, a $b$-matching $F$ is said to be a $\gamma$-approximate maximal matching if there exists a maximal matching with size $M$ such that $F$ has size at least $M/\gamma$. If there is a maximal $b$-matching of size $M$ where $F$ has size at least $M/\gamma$, then we say that $F$ is a $\gamma$-approximate $b$-matching. with probability at least $1-1/\mathrm{poly}(n)$. Then, when decoded from the billboard, with probability at least $1/2$, there is a node matched with at least $\Omega\left(\frac{\log(n)}{\varepsilon}\right)$ other nodes, regardless of the value of $\gamma>0$.
@@ 1.1.3 Node Privacy
In the node differential privacy setting, two graphs are considered neighbors if one can be obtained from the other by replacing all edges incident on any one node. It is a significantly more challenging setting than edge privacy. We show in Section 6 that the ideas we developed for edge-DP in general graphs can be combined with a “matching sparsifier” [solomon2021local] to give the first implicit node-DP algorithms for matching in general graphs. Specifically, we show the following upper bound results in node-DP within Section 6, where $\alpha$ is the arboricity [^3] Recall that a graph has arboricity $\alpha$ if it can be partitioned into $\alpha$ edge-disjoint forests. of the input graph.
Theorem 1.8 (Node-DP Approximate Maximum Matching).
Let $\eta\in(0,1]$, $\varepsilon\in(0,1)$, and $\alpha$ be the arboricity of the input graph. There is an $\varepsilon$-node DP algorithm that outputs an implicit $b$-matching. Moreover, with high probability, (i) $b=O\left(\frac{\alpha\log(n)}{\eta\varepsilon}+\frac{\log^{2}(n)}{\eta^{2}\varepsilon^{2}}\right)$, and (ii) the implicit solution has the size of the $(2+\eta)$-approximate maximum matching.
We can alternatively present our approximation guarantees in terms of $b$-matchings, which are detailed in Section 10.
Theorem 1.9 (Node-DP Approximate Maximum $b$-Matching).
Let $\eta\in(0,1]$, $\varepsilon\in(0,1)$, and $\alpha$ be the arboricity of the input graph. There is an $\varepsilon$-node DP algorithm that, with high probability, outputs an (implicit) $b$-matching in the billboard model, for $b\geq\frac{(1+\eta)^{2}}{1-\eta}b^{\prime}+\Omega\left(\frac{\alpha\log(n)}{\eta^{3}\varepsilon}+\frac{\log^{2}(n)}{\eta^{3}\varepsilon^{3}}+\frac{b^{\prime}\log(n)}{\eta^{2}\varepsilon}\right)$, that has the size of a $\left(2+\eta\right)$-approximate maximum $b^{\prime}$-matching.
Our techniques also lead to improved results for the setting of [hsu2014private]: bipartite graphs with node differential privacy in the billboard model. In this model, they view the nodes in the “left side” of the bipartition as goods and the “right side” nodes of the graph as bidders for these goods. Two such bipartite graphs are neighboring if one can be obtained from the other by adding/deleting any one bidder (and all its incident edges).
[hsu2014private] show guarantees of the following form. Suppose each item has supply $s$, meaning that each item can be matched with $s$ bidders (let us call these $s$-matchings); if $s=\Omega(\log^{3.5}{n}/\varepsilon)$, then they give an algorithm which (implicitly) outputs a $(1+\eta)$-maximum matching while guaranteeing differential privacy. [^4] They also consider the weighted version of the problem and further generalizations where each bidder has gross substitutes valuations. Our techniques also extend to these cases with improved results, but we limit our discussion here to matchings as that is the focus of this work. They also show that $s=\Omega(1/\sqrt{\eta})$ is necessary to obtain any guarantees, leaving open the question of tightening these bounds. We substantially tighten the upper bound in Section 7, showing that $s=\Theta(\log{n}/\varepsilon)$ suffices:
@@
If the supply is at least $s\geq\Omega\left(\frac{\log(n)}{\eta^{4}\varepsilon}\right)$, Algorithm 5 outputs a $(1+\eta)$-approximate maximum (one-sided) $s$-matching with high probability.